A restaurant offered cooking classes on 12 out of the 30 days in November. What decimal is equivalent to

the fraction of days in November that classes were offered at the restaurant?


1
Expert's answer
2021-09-14T06:05:03-0400
"\dfrac{12}{30}=\dfrac{2(6)}{5(6)}=\dfrac{2}{5}"

"\dfrac{2}{5}"